OpenStudy (anonymous):

Last year, Amy invested her money in two purchases. She purchased a certificate of deposit for 12,000 that paid 6% interest per year and purchased 4000 in corporate bonds paying 5% interest per year. Do not do any rounding What was the total interest earned at the end of 1 year? What was the percent interest for her total investment.

OpenStudy (gorv):

interest= 12000*6/100+4000*5/100 =120*6+40*5 =720+200=920

OpenStudy (gorv):

total invest=12000+4000=16000 %interest= (920/16000)*100=92/16=5.78 %

OpenStudy (gorv):

@AwesomeAries

OpenStudy (anonymous):

thank you!

OpenStudy (gorv):

wlcme :)